Tech Staff Meeting Notes April 2, 2009

The top priority is always repair of down stations.

The task list updated February 6, 2009 needs to be updated again. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


Ft Payne station
SM trigger upgrades
SM sample rate fix

News & Highlights

James got an extension to the end of the calendar year. He and Chris will work together to schedule one or a couple Davis Brain Dump seminars to prepare for the GPS transition. The "analog" parts will become the responsibility of Chris et al. and the "digital" parts will pass to Steve and Al. Data transfer, processing, etc will continue to be done by Bob Smalley and John Paul.

The search to fill the open research equipment tech I position was not successful. We will conduct another search at a future date.

ARRA funding (American Recovery and Reinvestment Act of 2009, aka stimulus funding) is forthcoming. Our priorities will be yr3 hardware for all SM stations, upgrading NMSZ 40T/dm24 stations to trillium/episensor/datalogger (data logger will likely be reftek), and upgrading all ET stations to digital (probably reftek) with possible addition of episensors (they will stay SP stations with S-13 or L4).

NetOpsIII will be in Menlo Park on May 4-5, 2009. It will be primarily focused on transitioning free field NSMP stations to RSN's.

Horton and Mitch are working with the AR Geological Survey to install a 6 station AR broadband network. Haydar also got earmark funding and this may be combined with the AGS funding. Adopting a network of EarthScope TA stations (Transportable Array) could be a good option.


Short Period & Broadbands

Holly's list:

Down or unusable

Isis Oxidation problem

Gain Channel problem

PWLA There is still significant telemetry interference that seems to be greatest during the work day.

Several of the stations can get improved telemetry by transmitting directly to the node. The original configuration changed enough so that there is room in the band. Jim will investigate.

NODE Tasks

MKTA: ISIS telemetry upgrade, return backup AC units to doghouses, finish ISIS routine service

LNXT: Replace XLAK clock and tie generator status alarms into lnxtworm

NMAD: Repair barbed wire fence damaged by falling ice

AZTN: Upgrade PC and replace XLAK clock

Steve has contacted telecom to order dsl for the three NMSZ nodes and the new Boswell node. Need to gather up a contact list (e.g. everyone's home phone, power companies, etc) and put it on webworm (password protected).


The gcf2ew processes on shakeworm are now replaced with scream/scream2ew on rhydra1. CVTN & ASAR still on gcf2ew pending the switch in the firewall to the rhydra ip:port.

SEAR was moved back to gcf2ew because it doesn't reconnect after an outage for some unknown reason.

CVTN optical isolation. Need to upgrade to year 3 hardware.

RDST needs upgrade to year 3 hardware with router.

NAIT: Couldn't get the NWS cable modem to cooperate at NAIT (firewall?). Dean was going to ask Comcast about it and maybe get a new modem. Steve will follow up.

Need to complete firmware upgrade on ASAR, SEAR, GILT, and CVVA.

Now that CMG5 triggers actually contain data, we need to tune the trigger level. Mitch sent explanatory info and guidelines to Al and Steve to tune.

Most of the strong motion stations can run at 100 sps without any modifications. We will install a router on the few that have problems (RDST will be the test case).

Remove and repair recorders 1 and 6, Repair or remove and repair recorder 2. The recorder disks should be reformatted as a routine maintenance task next time access is available. The constant read/write to the fat file systems lead to fragmentation. Don't ya just love dos? Wireless link to # 5 is unstable. Wireless link to #7 has been down for several weeks, not sure of recorder status. Verify if recorders will work with windowsXP while we have several here at CERI.

Drilling is complete and final cleanup is done. Wilson will evacuate water from 100' hole and use downhole camera to try and find leak.

We have the digitizer panel & one of the borehole instruments from VLF. Chris has the PVC enclosure, cabinet & vault ready to install. TDOT will allow CERI to install an earthworm PC at the Boswell TMC. CERI will need DSL service at the TMC. This eliminates an additional wireless link from Boswell to Richardson.

Mud Island vertical array is operational. Not really, functional is a better word until we can get it online to qc/vet data.


Next trip scheduled for April 27.


RCGA charging circuitry at is old and primitive and in need of an upgrade.

CCRT - Needs a rebuild after a bad lightning strike.

GMG - Need battery box solar reg mounting plate and need to fuse the battery.


GFM - Fixed the bad channel but still needs some work to repair storm damage. The NC state park service is purchasing GFM. We already have a couple permissions from them and John initiated the paperwork for the GFM permission.

TRYN: the forest service wants to remove the tower. There will be a new 180' tower and we have permission to be on it. We also have permission to be in the building (rack mount only). Can we convert this station to AC?

SWET: This is on hold pending ARRA projects. We need a plan for adding a strong motion instrument; a cmg5 will fit in the vault. AC power will be expensive and difficult. We need to nail down the hardware so we can figure out how much solar we need. Should we switch to ethernet radios (currently have serial) and have a dual-port terminal server? Should we use the intuicom radios they use at UUSS? Can there be simplification and hardening of the current path through the bell tower? Can we move the PC in Bran's office to someplace more secure and inaccessible?

CPCT: The aluminum enclosure with the GPS breakout board at CPCT is corroded, fragile, and needs replaced. Need to re-work lighting and pest chaser. John has a pest chaser and will use some landscape or pool hardware for the lighting. A new breakout and transient suppression enclosure is ready. Will CPCT also get a strong motion upgrade? Needs a locking tower box.

John found a good host for a Ft Payne broadband. There is AC (or rather the host is willing to run power). The ISP is also willing to provide free bandwidth though this may not last long. There is a second possibility where we could telemeter to Jacksonville State (

There is SOH info superimposed on the pII gain channel. Jim and John will get the codes to Mitch who will investigate the feasibility of decoding software. Mitch looks at this for a day. It can be done but won't be easy to do robustly. Some of the documentation appears contradictory and it isn't clear which bit is what. This is not needed if we remove the pII and install refteks.


PTGV is back online. The wireless router in the Delta center offices was repackaged and reinstalled now the field site needs an equipment enclosure upgrade. Done and Bob is very pleased with the quality of the work.

Camp Robinson doesn't answer calls. Might be able to combine with a visit to SEAR.

Need a routine PM visit to all stations.

Davis brain dump scheduled for April 9.


Need to do a CMG6TD huddle test.

JerMing wants his class to do a two week array deployment.


Leaky seals on the giddings are repaired.

What to do with the Steiner brain dump DVD's?

How do we make John's database accessible from a central server by everyone authorized (and not by anyone not authorized)? One possibility is to put it in seisadm on smeagol accessible via gaia and let John remote desktop to a seisnet xp server that can access it (everyone else within the ceri domain can access directly via gaia).

Steve and Al are setting up a feed of Oklahoma data to be forwarded to the NEIC. The scream2scream connection didn't seem to be all that stable and OKGS could use a better way to archive events before they leave the buffer. Steve will help them get ew going and then we'll use the ew export. We need to make sure to add relevant channels to the NEIC export when available and contact Dan McManus dmcmanus) at USGS Golden when they're going.

Rotational sensor is feeding an earthworm but still has IRIG crosstalk problem to fix.

Inventory. Apparently other networks are resisting and INV is only fully functional at NEIC. The dataless seed generation works well if you already have the ascii dataless and not at all if you don't. We might use datascope for this instead (talk about full circle--withers started that in 1997).

John has developed a working access database that should work for tracking station information and maintenance.

The new Suburban is in service and Dave is working on customizing for field equipment, tools etc. Mac needs to decide which truck he wishes to decommission (old surburban 148 or the blue pickup). Its a painful thing to watch--its as if we asked him to sacrifice a child. QUIT STALLING MAC AND GET THIS DONE PRONTO!!

Pink Palace is working with Gary and Michelle on a new perc-like display area. We'll pull the receiver etc, when they start pulling out the old display (or sometime earlier if an opportunity arises).

The Transportable Array (TA) component of the U.S. Array observatory of the Earthscope Facility will be installed in the NMSZ in 2011 and permissioning in AR and MO will be the summer of 2009. Jim and Chris made contact lists and Mitch passed them on.

(Old note from January) Mitch briefly reviewed the desired 106x6 network of desirable TA stations to adopt. These are the 106 TA stations within the ANSS RSN footprint for the NSMZ and WVSZ and the ETSZ and CSZ. The "by 6" implies three components for broadband and three of strongmotion. Considerable funds must be secured for the capitalization costs and for the long term operation and maintenance.


3876 Power The power for the new lighting on the east side of the general parking lot to the west of CERI is coming from 3876. Chris will check it out to assess the added lightning hazard and see if anything can be done about it. Might be hard to do without the wiring diagram. Chris will make one if needed as time permits.

3890 Power Chris is going to make a diagram for 3890. Good luck son.